OpenStudy (anonymous):

is the following sequence arithmetic? 20, 24, 28, 32,... a. yes b. no

OpenStudy (anonymous):

32 - 28 = 4 28 - 24 = 4 24 - 20 = 4 See a pattern?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thankyou kdwndr:)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

and thankyou meverette04!!!